Today we investigated the issue experienced yesterday with the ID of the WE.
We performed a measument in order to recompute the calibration for the accelerometeres. We found out that ACC_H1 was the one miscalibrated, but the correct compensator was the one used before the last update. One possibility is that the accelerometer switches between two states. The reason of this behaviour is not clear. Some investigation is needed.
In the meantime, we reloaded the older compensation filter, and we modified the DSP card adding one switch for the two compensators in order to have the possibility to move between one 'state' and the other.